OpenStudy (anonymous):

Use the Distributive Property to find the price of 4 CDs that cost $14.99 each.

OpenStudy (anonymous):

@Directrix is the answer $59.96

Directrix (directrix):

Well, 4* $14.99 = $59.96 but I don't know what the question is.

OpenStudy (anonymous):

14.99 * 4 = 59.96

Directrix (directrix):

Use the Distributive Property

OpenStudy (anonymous):

yeah i agree with @Directrix Sir :)

OpenStudy (anonymous):

14.99 + 14.99 + 14.99 + 14.99 =

OpenStudy (skullpatrol):

Try to use the distributive property as the question asks.

OpenStudy (anonymous):

14.99 + 14.99 + 14.99 + 14.99 = 59.96

Directrix (directrix):

4* ($14 +.99) = 4* 14 + ?*?

OpenStudy (anonymous):

Distributive Property a(b + c) = ab + ac

Directrix (directrix):

The Distributive Property is handy to help you get rid of parentheses or in this case, to create them. a(b + c) = ab + ac where a, b, and c are real numbers.